Question 9 Consider the following two stocks: Stock A Stock B Expected Return Variance 1196 696 0.0324 0.0196 If the correlation of their return is equal to 0.3, what is the standard deviation of a portfolio of the two stocks whose weight in stock A is equal to 0.3? (Answers rounded to two decimal digits)

hi guys how can i solve this with good explanation
Answer #1

Standard deviation of stock A, d1 = (0.0324)1/2 = 0.18 or 18%

Standard deviation of stock B, d2 = (0.0196)1/2 = 0.14 or 14%

Weight of stock A = 0.3

Weight of stock B = 1 - 0.3 = 0.7

Standard deviation of portfolio = ( W12d12 + W22d22 + 2 * W1 * W2 * correlation * standard deviation1 * standard deviation2)1/2

Standard deviation of portfolio = ( 0.320.182 + 0.720.142 + 2 * 0.3 * 0.7 * 0.3 * 0.18 * 0.14)1/2

Standard deviation of portfolio = ( 0.09*0.0324 + 0.49*0.0196 + 0.003175)1/2

Standard deviation of portfolio = ( 0.002916 + 0.009604 + 0.003175)1/2

Standard deviation of portfolio = 0.1253 or 12.53%